# Wav-KAN: Wavelet Kolmogorov-Arnold Networks

### v1

@@ Abstract
In this paper  [^1] Another title for the paper can be ”Wavelet for Everybody”., we introduce Wav-KAN, an innovative neural network architecture that leverages the Wavelet Kolmogorov-Arnold Networks (Wav-KAN) framework to enhance interpretability and performance. Traditional multilayer perceptrons (MLPs) and even recent advancements like Spl-KAN [1] face challenges related to interpretability, training speed, robustness, computational efficiency, and performance. Wav-KAN addresses these limitations by incorporating wavelet functions into the Kolmogorov-Arnold network structure, enabling the network to capture both high-frequency and low-frequency components of the input data efficiently. Wavelet-based approximations employ orthogonal or semi-orthogonal basis and also maintains a balance between accurately representing the underlying data structure and avoiding overfitting to the noise. Analogous to how water conforms to the shape of its container, Wav-KAN adapts to the data structure, resulting in enhanced accuracy, faster training speeds, and increased robustness compared to Spl-KAN and MLPs. Our results highlight the potential of Wav-KAN as a powerful tool for developing interpretable and high-performance neural networks, with applications spanning various fields. This work sets the stage for further exploration and implementation of Wav-KAN in frameworks such as PyTorch, TensorFlow, and also it makes wavelet in KAN in wide-spread usage like nowadays activation functions like ReLU, sigmoid in universal approximation theory (UAT).
@@ I Introduction
Incorporating Wav-KAN in neural networks makes increasingly explainable components that can achieve similar state-of-the-art performance across diverse tasks  [^3] Some of the proofs and experiments will be presented in the final version of the paper.
Section II discusses KAN and its generalization to multi-layers KAN with wavelet as function approximation (i.e., activation function). Section III presents Continuous Wavelet Transform, especially the definition and criteria for being a mother wavelet which can be used as basis for function approximation. The comparison of Wav-KAN, Spl-KAN and MLPs is given in IV. Some experiment results will be given in Section V. Finally, Section VI concludes the paper.
@@ II-B From Theory to Networks
In summary, traditional MLPs use fixed nonlinear activation functions at each node and linear weights (and biases) to transform inputs through layers. The output at each layer is computed by a linear transformation followed by a fixed activation function. During backpropagation, gradients of the loss function with respect to weights and biases are calculated to update the model parameters. In contrast, KANs replace linear weights with learnable univariate functions placed on edges rather than nodes. In the nodes, we just have a summation of some univariate functions from previous layers. Each function is adaptable, allowing the network to learn both the activation and transformation of the inputs. This change leads to improved accuracy and interpretability, as KANs can better approximate functions with fewer parameters and avoid the curse of dimensionality. During backpropagation in KANs, the gradients are computed with respect to these univariate functions, updating them to minimize the loss function. This results in more efficient learning for complex and high-dimensional functions.
@@ III Continuous Wavelet Transform
IV Wav-KAN or Spl-KAN or MLPs?
@@
Moreover, Spl-KAN heavily depends on the grid spaces, and for better performance, it requires increasing the number of grids  [^6] Which is equivalent to decrease the spaces between the grids. This helps Spl-KAN accuracy to be improved, though, it brings two disadvantages. First it needs curvefitting which is a cumbersome and computational expensive operation, and also while we increase the number of grids, loss has some jumps  [^7] Loss will be increased for some training, then, it works better. See Figure 2.3 [1] for better illustration.. Fortunately, wavelet is safe from such computations and deficiencies.
@@
V Simulation Results
In this section, we present the results of our experiments conducted using the KAN (Kernel-based Artificial Neural network) model with various wavelet transformations on the MNIST dataset, utilizing a training set of 60,000 images and a test set of 10,000 images. It is important to note that our objective was not to optimize the parameters to their best possible values, but rather to demonstrate that Wav-KAN performs well in terms of overall performance. We have incorporated batch normalization into both Spl-KAN and Wav-KAN, resulting in improved performance. The wavelet types considered in our study include Mexican hat, Morlet, Derivative of Gaussian (DOG), and Shannon (see Table II). For each wavelet type and also Spl-KAN, we performed five trials, training the model for 50 epochs per trial.
@@ V Simulation Results
Both Wav-KAN and Spl-KAN have the structure (number of nodes) of [28*28,32,10] 999[first layer nodes, middle layer nodes, output nodes]. Although we enhanced Spl-KAN by using spline order 3 and a grid size of 5, this approach is computationally much more expensive compared to Wav-KAN. We employed AdamW optimizer [21, 22], with learning rate of 0.001 with weight decay of $10^{-4}$. Loss is cross entropy.
@@
Figures 2 and 3 show the result of training accuracy and test accuracy of Wav-KAN in comparison to Spl-KAN. To not clutter up, we just show the result of Derivative of Gaussian (DOG) and Mexican hat wavelet. These have been shown as a sample. By fine tuning, and using all the freedom of wavelet (like frequency of sinusoid and variance of Gaussian), Wavelet shows significant superiority  [^10] We have done a lot of such experiments and we will publish them soon.. While Spl-KAN has better performance in training, which is because of overfitting to data, a lot of wavelet types have shown superior performance with respect to Spl-KAN. Indeed, for this experiment we set the variance of the Gaussian to be 1; though, we can find better variances by grid search or making it a learnable parameter. Wavelet makes a balance by not fitting to the noise in the data.

### v2

@@ Abstract
In this paper  [^1] Another title for the paper can be ”Wavelet for Everybody”., we introduce Wav-KAN  [^2] Wav-KAN refers to a family of continuous and discrete wavelet transforms in KAN., an innovative neural network architecture that leverages the Wavelet Kolmogorov-Arnold Networks (Wav-KAN) framework to enhance interpretability and performance. Traditional multilayer perceptrons (MLPs) and even recent advancements like Spl-KAN [1] face challenges related to interpretability, training speed, robustness, computational efficiency, and performance. Wav-KAN addresses these limitations by incorporating wavelet functions into the Kolmogorov-Arnold network structure, enabling the network to capture both high-frequency and low-frequency components of the input data efficiently. Wavelet-based approximations employ orthogonal or semi-orthogonal basis and also maintains a balance between accurately representing the underlying data structure and avoiding overfitting to the noise. While continuous wavelet transform (CWT) has a lot of potentials, we also employed discrete wavelet transform (CWT) for multiresolution analysis which obviated the need for recalculation of the previous steps in finding the details  [^3] We demonstrate its efficacy in capturing detailed signal characteristic. and efficiently combines local detailed information where the data points are dense with broader trends where the data points are sparse. Analogous to how water conforms to the shape of its container, Wav-KAN adapts to the data structure, resulting in enhanced accuracy, faster training speeds, and increased robustness compared to Spl-KAN and MLPs. Our results highlight the potential of Wav-KAN as a powerful tool for developing interpretable and high-performance neural networks, with applications spanning various fields. This work sets the stage for further exploration and implementation of Wav-KAN in frameworks such as PyTorch, TensorFlow, and also it makes wavelet in KAN in wide-spread usage like nowadays activation functions like ReLU and sigmoid in universal approximation theory (UAT). The codes to replicate the simulations are available at https://github.com/zavareh1/Wav-KAN.
@@ I Introduction
Incorporating Wav-KAN in neural networks makes increasingly explainable components that can achieve similar state-of-the-art performance across diverse tasks  [^5] Some of the proofs and experiments will be presented in the final version of the paper..
Section II discusses KAN and its generalization to multi-layers KAN with wavelet as function approximation (i.e., activation function). Section III presents Continuous Wavelet Transform, especially the definition and criteria for being a mother wavelet which can be used as basis for function approximation. The comparison of Wav-KAN, Spl-KAN and MLPs is given in V. Some experiment results will be given in Section VI. Finally, Section VII concludes the paper.
@@ II-B From Theory to Networks
In summary, traditional MLPs use fixed nonlinear activation functions at each node and linear weights (and biases) to transform inputs through layers. The output at each layer is computed by a linear transformation followed by a fixed activation function. During backpropagation, gradients of the loss function with respect to weights and biases are calculated to update the model parameters. In contrast, KANs replace linear weights with learnable univariate functions placed on edges rather than nodes. In the nodes, we just have a summation of some univariate functions from previous layers. Each function is adaptable, allowing the network to learn both the activation and transformation of the inputs. This change leads to improved accuracy and interpretability, as KANs can better approximate functions with fewer parameters. During backpropagation in KANs, the gradients are computed with respect to these univariate functions, updating them to minimize the loss function. This results in more efficient learning for complex and high-dimensional functions.
@@ III Continuous Wavelet Transform
IV Discrete Wavelet Transform
The Discrete Wavelet Transform (DWT) is a widely used method in signal processing that decomposes a signal into different frequency components with high efficiency [14]. Unlike the Continuous Wavelet Transform (CWT), which analyzes the signal at every possible scale and translation, the DWT uses discrete sampling intervals which is suitable for digital applications. When the data points are close together (high sampling rate), the wavelet basis can zoom in to capture the high-frequency details (fine details) in that local area. When the data points are spread out (low sampling rate), the wavelet basis can zoom out to capture the low-frequency trends (overall shape) using global information. We see some examples uniform and irregular samples of data. Indeed, by employing wavelet we combine local detailed information where the data points are dense with broader trends where the data points are sparse.
The DWT employs a set of base functions derived from a single mother wavelet through scaling and translation. These base functions are orthonormal, ensuring that the DWT provides a compact and non-redundant representation of the signal. By applying high-pass and low-pass filters iteratively to the signal, the DWT produces approximation and detail coefficients at various levels of resolution, allowing for a multi-resolution analysis.
[…]
$a_{j}(k)=\sum_{n}g(n)\phi_{j,k}(n)$ (16)
$d_{j}(k)=\sum_{n}g(n)\psi_{j,k}(n)$ (17)
where:
  • $g(n)$ is the discrete signal.
  • $\phi_{j,k}(n)$ is the scaling function at scale $j$ and position $k$.
  • $\psi_{j,k}(n)$ is the wavelet function at scale $j$ and position $k$.
  • $a_{j}(k)$ represents the approximation coefficients.
  • $d_{j}(k)$ represents the detail coefficients.
The original signal can be reconstructed from its wavelet coefficients using the inverse DWT, ensuring that no information is lost during the transformation process. This reconstruction is given by:
$g(n)=\sum_{k}a_{j}(k)\phi_{j,k}(n)+\sum_{j=1}^{J}\sum_{k}d_{j}(k)\psi_{j,k}(n)$ (18)
where $J$ is the number of decomposition levels.
The DWT’s ability to provide both time and frequency localization, coupled with its computational efficiency, makes it an indispensable tool in various fields such as image compression, noise reduction, and feature extraction.
Multi-resolution analysis (MRA) using wavelets, specifically discrete wavelet transforms (DWT), is a powerful technique in signal processing and data analysis. MRA allows the decomposition of a signal into different levels of detail by using wavelets, which are localized functions. The DWT achieves this by recursively applying high-pass and low-pass filters to the data/signal, creating approximations and details at various resolutions. This process provides a hierarchical framework, where each level captures different frequency components of the signal, making it easier to analyze localized features and transient phenomena. The ability to zoom in on fine details while retaining the broader structure of the signal is a key advantage of MRA, making it useful in applications such as image compression, noise reduction, and feature extraction.
In the context of discrete wavelet transforms, the data/signal is represented as a sum of wavelet functions at different scales and positions. In wavelet analysis, once we compute the initial set of coefficients, we do not need to recompute them when calculating additional coefficients at finer resolutions. This efficiency arises because wavelet transforms leverage previously computed coefficients to generate further details. Consequently, the hierarchical nature of the multi-resolution analysis ensures that the earlier computations remain valid and reusable, streamlining the process and significantly reducing the computational overhead. This property makes wavelet-based methods particularly advantageous for iterative and real-time signal processing applications. The multi-resolution property of wavelets is particularly beneficial for analyzing non-stationary signals, which have frequency components that vary over time. By breaking down the signal into different resolution levels, MRA enables a more comprehensive understanding of its structure and characteristics. Moreover, the DWT’s ability to provide both time and frequency localization offers a significant advantage over traditional Fourier transforms, which only provide frequency information  [^8] We will add some simulations to this section..
In summary, wav-kan by employing wavelet adapts to the varying density of data points, using more detailed information where there are more data points and less detailed information where there are fewer data points. This approach is particularly useful for real-world data that often suffers from irregular sampling and data dropouts.
V Wav-KAN or Spl-KAN or MLPs?
@@
Moreover, Spl-KAN heavily depends on the grid spaces, and for better performance, it requires increasing the number of grids  [^10] Which is equivalent to decrease the spaces between the grids. This helps Spl-KAN accuracy to be improved., though, it brings two disadvantages. First it needs curvefitting which is a cumbersome and computational expensive operation, and also while we increase the number of grids, loss has some jumps  [^11] Loss will be increased for some training, then, it works better. See Figure 2.3 [1] for better illustration.. Fortunately, wavelet is safe from such computations and deficiencies and if one wants to capture more details, DWT efficiently does that without recalculation of the previous steps.
@@
VI Simulation Results
In this section, we present the results of our experiments conducted using the KAN (Kernel-based Artificial Neural network) model with various continuous wavelet transformations (CWTs)  [^13] We add more simulations including some discrete wavelet ones for MRA, etc. on the MNIST dataset, utilizing a training set of 60,000 images and a test set of 10,000 images. It is important to note that our objective was not to optimize the parameters to their best possible values, but rather to demonstrate that Wav-KAN performs well in terms of overall performance. We have incorporated batch normalization into both Spl-KAN and Wav-KAN, resulting in improved performance. The wavelet types considered in our study include Mexican hat, Morlet, Derivative of Gaussian (DOG), and Shannon (see Table II). For each wavelet type and also Spl-KAN, we performed five trials, training the model for 50 epochs per trial.
@@ VI Simulation Results
Both Wav-KAN and Spl-KAN have the structure (number of nodes) of [28*28,32,10] 141414[first layer nodes, middle layer nodes, output nodes]. Although we enhanced Spl-KAN by using spline order 3 and a grid size of 5, this approach is computationally much more expensive compared to Wav-KAN. We employed AdamW optimizer [21, 22], with learning rate of 0.001 with weight decay of $10^{-4}$. Loss is cross entropy.
@@
Figures 2 and 3 show the result of training accuracy and test accuracy of Wav-KAN in comparison to Spl-KAN. To not clutter up, we just show the result of Derivative of Gaussian (DOG) and Mexican hat wavelet. These have been shown as a sample. By fine tuning, and using all the freedom of wavelet (like frequency of sinusoid and variance of Gaussian), Wavelet shows significant superiority  [^15] We have done a lot of such experiments and we will publish them soon.. While Spl-KAN has better performance in training, which is because of overfitting to data, a lot of wavelet types have shown superior performance with respect to Spl-KAN. Indeed, for this experiment we set the variance of the Gaussian in wavelets to be 1; though, we can find better variances by grid search or making it a learnable parameter. Also, one may like to consider $\omega_{0}$ in Morlet and parameters of frequency and/or windowing in Shannon wavelet learnable. Indeed, we can use different degrees of freedom that we have to make the shape of wavelets more flexible. Wavelet makes a balance by not fitting to the noise in the data.